Falconmaster:

What is an equation of the line that passes through (0, 8) and (4, 0)? y = -2x + 4 y = 2x + 4 y = -2x + 8 y = 2x + 8

Falconmaster:

@AMQ

Bob:

www.desmos.com

AMQ:

C

AMQ:

100% sure :p

iGreen:

Do you want to learn how to do it or cheat and fail math the rest of your life?

AMQ:

it's ok to rage when you don't get the medals green. there there..... *pats head*

Bob:

I thought direct answers were illegal on this site...

Falconmaster:

nope

AMQ:

igreen with his paragraph of bull shi't VVVVVVVVVVVv

AMQ:

trust me on this ;)

iGreen:

To figure out the problem, find the slope of the two points. (0, 8), (4, 0) x1 y1 x2 y2 \(\sf \dfrac{y_2 - y_1}{x_2-x_1}\) After you find the slope you plug it in with one of the points into point slope form. \(\sf y - y_1 = m(x - x_1)\) Where \(\sf m\) is the slope and \(\sf (x_1, y_1)\) is a point on the line. Then you can convert it to slope-intercept form.

AMQ:

(;